OpenStudy (anonymous):

Solve for x. Round to nearest hundredth. 2=ln(x+7)

Vocaloid (vocaloid):

rewriting both sides in base e gives us e^2 = x + 7 solve for x

OpenStudy (anonymous):

What is e?

OpenStudy (solomonzelman):

e is the base of the natural log.

OpenStudy (solomonzelman):

In other words, \(\ln(a)\) is same as \(\log_e(a)\)

OpenStudy (anonymous):

So how do i find x?

OpenStudy (solomonzelman):

Use: \(\log_a{b}=c ~~~~~~\Longrightarrow ~~~~~~a^c=b\)

OpenStudy (solomonzelman):

this is the rule./

OpenStudy (solomonzelman):

\(\log_e(x+7)=2\) \(e^2=x+7\) \(x=?\)

OpenStudy (anonymous):

I understand the process but i dont understand how to find x

OpenStudy (solomonzelman):

you have just a single step left.

OpenStudy (anonymous):

Would it be 2.64 ??

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

(your answer will involve e^2)

OpenStudy (anonymous):

I have no idea.m

OpenStudy (solomonzelman):

subtract 7 from both sides

OpenStudy (anonymous):

-7?

OpenStudy (solomonzelman):

e^2=x+7 subtract 7 from both sides and you get?

OpenStudy (anonymous):

E^2-7 =x

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

Would the answer just be -7? It has to be numbers..i cant use letters

OpenStudy (solomonzelman):

approximtae e^2

OpenStudy (solomonzelman):

http://www.wolframalpha.com/input/?i=e%5E2-7

OpenStudy (solomonzelman):

about 0.3891

OpenStudy (anonymous):

Since it says to round to the nearest hundredth would it just be 0.389

OpenStudy (anonymous):

??

OpenStudy (solomonzelman):

hunderedth is the 2dn decimal place

OpenStudy (solomonzelman):

2nd **

OpenStudy (anonymous):

Thank you for all your help!!